OpenStudy (seanm95):

Consider an object moving along the curve y=sin(x), starting at the origin. Assuming that the object's x-coordinate, s, is increasing at a rate of 3 ft/sec, what is the rate of change of p at the moment when s=pi/4? (P denotes the x-intercept of the line tangent to the curve y=sin(x) at the point (s, sin(s))). Is p at that moment moving towards or away from the origin?

OpenStudy (phi):

do you have the equation of the line tangent ?

OpenStudy (seanm95):

There is no equation given.

OpenStudy (peachpi):

y = sin (x) dy/dx = cos (x) tangent line at (s, sin(s)) y - sin (s) = dy/dx (x - s) y - sin (s) = cos (x) (x - s) x-intercept is at (p, 0): 0 - sin (s) = cos (p)(p - s) -sin (s) = cos (p) (p - s) then take the derivative to solve for dp/dt -cos(s) ds/dt = cos (p) (dp/dt - ds/dt) - (p - s)sin (p)(dp/dt) I think...

OpenStudy (irishboy123):

@peachpi you're kinda kinda where i come out but i can't totally reconcile our separate approaches we have position vector \(\vec r = <s, \sin s>\) and so velocity vector \(\vec r' = \vec v = <\dot s, \cos s ~ \dot s>\) The velocity vector is the tangent vector so the tangent line is parameterised as \(\vec l = <s, \sin s> + \lambda <\dot s, \cos s ~ \dot s>\) This will cross the x axis at \(y = \sin s + \lambda \cos s \dot s = 0\) \(\implies \lambda = - \dfrac{\sin s}{\cos s ~ \dot s} = - \dfrac{\tan s}{ \dot s}\) so x value, ie p, is \(p = s + \lambda \dot s = s - \tan s\) At that point \(\dot p = \dot s - \sec^2 s ~ \dot s\) \(\dot p = - \dot s \tan^2 s\) reality check....does this make sense at \(s = 0, \frac{\pi}{2}\) etc. and it kinda does so plug in \(s = \frac{\pi}{4}, \dot s = 3\)